OpenStudy (anonymous):

what is the second derviteve of f(x)=1-8a^-3

OpenStudy (anonymous):

did you find the first derivative?

OpenStudy (anonymous):

the first derivative is f'(x) = -3*-8a^-3-1 the second derivative is f''(x) = (-3-1)(-3*-8)a^-3-1-1

Parth (parthkohli):

Uh, shouldn't it be \(1 - 8x^{-3}\)?

OpenStudy (anonymous):

i assume it is x not a. \[f \left( x \right)=1-8x ^{-3}\] \[f \prime \left( x \right)=0-8\left( -3 \right)x ^{-4}=24x ^{-4}\] \[f \prime \prime \left( x \right)=24\left( -4 \right)x ^{-5}=-96x ^{-5}\]

OpenStudy (joannablackwelder):

If it is a, then a is a constant and the derivative and second derivative are 0.